# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 1.1.2. Some signposts for Brownian Boost results
Our first, main task is a rigorous analysis of $\rho$-Brownian Boost. We will circumvent the characteristic challenges of instantaneous feedback loops for stochastic differential games by approximating the game with a discrete version played on a fine-mesh copy of the integers. For each $\rho\in(0,1]$, the time-homogeneous Markov-perfect Nash equilibria will be classified as a one-dimensional space, invariant under real shifts, and indexed by a ‘battlefield’ value in $\mathbb{R}$, with players at a given equilibrium staking intensely in a bounded region about the battlefield value.
When $\rho\in(0,1]$, the equilibria are described by a soon-stated $\rho$-parameterised ODE pair (2) that we will solve explicitly. The gameplay $X:[0,\infty)\to\mathbb{R}$ for the zero-indexed equilibrium solves the SDE
${\rm d}X_{u}=R_{\rho}(u)\,{\rm d}u+{\rm d}W_{u}\,,\,\,\,\textrm{with $W_{u}$ standard Brownian motion}\,$, (1)
whose drift term $R_{\rho}(u)$ equals $\frac{1-J(u)}{1+J(u)}$ where $J$ solves the ODE $\frac{{\rm d}J(u)}{{\rm d}u}\,=\,-8\rho^{2}\frac{J(u)^{2}}{(1+J(u))^{2}}$ with $J(0)=1$. Players fight hard for control as the counter passes close to the origin; as it drifts away according to the asymptotics $R_{\rho}(u)\to\pm 1$ seen in the respective limits $u\to\pm\infty$, the game enters a long low-stakes phase which typically reinforces the dominance of the leading player.
Now we present the ODE system, including its solutions and some important properties, that governs our characterization of $\mathrm{BB}(\rho)$ equilibria. Although we defer a presentation of the precise framework for strategies and gameplay in Brownian Boost, a basic aspect is needed to interpret the solutions we present. In playing $\mathrm{BB}(\rho)$, a player may in principle draw on a broad range of strategies determined by game history in choosing her stakes. We will restrict attention to a narrower class that includes all viable options according to an intuitively appealing principle akin to the Markov property: in the history of gameplay until a given moment, the one piece of data that should be determinative for deciding the stake rate is the present counter position $X_{t}$. Stake pairs that meet this condition are time-homogeneous and Markov perfect, and in shorthand we will call them time-invariant. By focussing on such pairs, we reinterpret the stakes specified in Section 1.1 as profiles $a,b:\mathbb{R}\to[0,\infty)$, with $a(x)$ and $b(x)$ denoting the rate at which Maxine and Mina stake at any moment $t\geq 0$ for which $X_{t}=x$. The drift $R_{\rho}(u)$ in the gameplay SDE (1) thus equals $\frac{a(X_{t})^{\rho}-b(X_{t})^{\rho}}{a(X_{t})^{\rho}+b(X_{t})^{\rho}}$.
Here is the ODE pair that will be shown to govern time-invariant Nash equilibria in $\mathrm{BB}(\rho)$. .
Definition 1.1.
$2\rho f^{1+\rho}g^{\rho}$ $=$ $\,\,\,\,\,\,\big(f^{2\rho}-g^{2\rho}\big)f\,+\,\tfrac{1}{2}f^{\prime}\big(f^{\rho}+g^{\rho}\big)^{2}\,$, (2)
A pair of functions $f,g:\mathbb{R}\to[0,\infty)$ is called default if $f(0)=1$ and $g(0)>0$.
The ODE pair arises as the coupled system of Hamilton-Jacobi-Bellmann [HJB] equations associated to the non-zero-sum game $\mathrm{BB}(\rho)$. In Section 5, we will explain this connection with a simple but non-rigorous argument. Using Markovian forward equations and stability under momentary perturbation of stake by a given player, the argument finds necessary conditions for a stake-profile pair $(a,b):\mathbb{R}\to[0,\infty)^{2}$ to be a Nash equilibrium. Associated to $(a,b)$ are $m,n:\mathbb{R}\to[0,\infty)$, the players’ mean total receipts as a function of initial counter location. Supposing differentiability, we have $m^{\prime}>0$ and $n^{\prime}<0$, since Maxine plays right and Mina left. The obtained conditions are that $(f,g)=(m^{\prime},-n^{\prime})$ is a $\rho$-Brownian Boost ODE pair. (In the theory of stochastic differential games, formal derivations of HJB equations would suppose sufficient differentiability; but, in contrast to $\mathrm{BB}(\rho)$, value functions often do not enjoy that regularity, and are instead exhibited rigorously as viscosity solutions [11]: see [20] and [5] respectively for zero- and non-zero-sum treatments.)
We will anaylse $\mathrm{BB}(\rho)$ rigorously by regularizing it as a discrete game in a suitable high-noise small-step limit; by doing so, we will substantiate (when $\rho\in(0,1]$) that the mentioned conditions characterise $\mathrm{BB}(\rho)$ equilibria. We defer explaining the discrete setup and how it scales to Brownian Boost for later in the introduction. For now, the prospect of such a characterization may provoke the question, how to solve the above pair of equations? We record the answer next, noting that currency revaluation permits us to consider only default solutions. Our analytic deductions hold whenever $\rho\in(0,\infty)$, even if the game-theoretic meaning of the $\mathrm{BB}(\rho)$ ODE pair is unsettled for $\rho>1$.
Definition 1.2.
For $\rho,x\in(0,\infty)$, let $S_{\rho}(x,\cdot):\mathbb{R}\to(0,\infty)$ denote the unique solution to the differential equation
$\frac{{\rm d}}{{\rm d}u}S_{\rho}(x,u)\,=\,-\frac{8\rho\,S_{\rho}(x,u)^{1+\rho}}{\big(1+S_{\rho}(x,u)^{\rho}\big)^{2}}\,,\qquad S_{\rho}(x,0)=x\,.$
Associate to this solution the pair of functions $f_{\rho}(x,\cdot),g_{\rho}(x,\cdot):\mathbb{R}\to(0,\infty)$ by means of
$f_{\rho}(x,r)\,=\,\exp\!\left\{2\int_{0}^{r}\Bigg(1-\frac{2}{\big(1+S_{\rho}(x,u)^{\rho}\big)^{2}}\,\Big(1+(1-\rho)S_{\rho}(x,u)^{\rho}\Big)\Bigg)\,{\rm d}u\right\}$
and
$g_{\rho}(x,r)\,=\,x\cdot\exp\!\left\{-2\int_{0}^{r}\Bigg(1-\frac{2}{\big(1+S_{\rho}(x,u)^{\rho}\big)^{2}}\,\Big((1-\rho)S_{\rho}(x,u)^{\rho}+S_{\rho}(x,u)^{2\rho}\Big)\Bigg)\,{\rm d}u\right\}$
for $r\in\mathbb{R}$. When $r<0$, the integrals are specified in the usual way: $\int_{a}^{b}h=-\int_{b}^{a}h$ for $a>b$.
Theorem 1.3.
Let $\rho\in(0,\infty)$. The space of default solutions to the system (2) is equal to
$\Big\{\big(f_{\rho}(x,\cdot),g_{\rho}(x,\cdot)\big):\mathbb{R}\to(0,\infty)^{2}\Big\}\,$,
where the index runs over $x\in(0,\infty)$. For each $x$, we have $g_{\rho}(x,\cdot)=f_{\rho}(x,\cdot)S_{\rho}(x,\cdot)$.
Given the pair $\big(f_{\rho}(x,\cdot),g_{\rho}(x,\cdot)\big)$, how to recover the stake profile pair $(a,b):\mathbb{R}\to[0,\infty)^{2}$ that is the putative associated Nash equilibrium; namely, for which $(m^{\prime},-n^{\prime})$ equals $\big(f_{\rho}(x,\cdot),g_{\rho}(x,\cdot)\big)$? The recipe is that $(a,b)$ equals $\big(a_{\rho}(x,\cdot,b_{\rho}(x,\cdot)\big)$ as now specified.
Definition 1.4.
For $x\in\mathbb{R}$, let $a_{\rho}(x,\cdot)$ and $b_{\rho}(x,\cdot)$ mapping $\mathbb{R}$ to $(0,\infty)$ be given by
$a_{\rho}(x,r)=2\rho\;\frac{f_{\rho}(x,r)^{\,1+\rho}\;g_{\rho}(x,r)^{\,\rho}}{\big(f_{\rho}(x,r)^{\rho}+g_{\rho}(x,r)^{\rho}\big)^{2}}$
and
$b_{\rho}(x,r)=2\rho\;\frac{f_{\rho}(x,r)^{\,\rho}\;g_{\rho}(x,r)^{\,1+\rho}}{\big(f_{\rho}(x,r)^{\rho}+g_{\rho}(x,r)^{\rho}\big)^{2}}\,.$
Definition 1.1 makes no reference to boundary conditions. The values $m_{\rho}(x,\infty):=\int_{-\infty}^{\infty}f_{\rho}(x,u)\,{\rm d}r$ and $n_{\rho}(x,-\infty):=\int_{-\infty}^{\infty}g_{\rho}(x,u)\,{\rm d}r$ are necessarily positive since we suppose $f$ and $g$ to be positive. As integrals of spatial derivatives for expected payoff, these quantities are the values of Maxine and Mina’s respective terminal rewards in the event of the given player’s victory.
Up to trivial symmetries, the pairs $\big(f_{\rho}(x,\bullet),g_{\rho}(x,\bullet)\big):\mathbb{R}\to(0,\infty)^{2}$ indexed by $x\in(0,\infty)$ specify all solutions of the $\rho$-Brownian Boost ODE pair. As we will discuss in Subsection 1.7.2, $m_{\rho}(x,\infty)$ and $n_{\rho}(x,-\infty)$ are equal for any given $x\in(0,\infty)$, so that essentially only one, symmetric, boundary condition is available.
This one-parameter family of solutions is in fact given (up to dilation) by a single solution and its translates, formed by replacing the domain variable $\bullet$ by $v+\bullet$ for some $v\in\mathbb{R}$. (See Proposition 5.4(2) for the relation between the variables $x$ and $v$.) The paradox of the existence of equilibria for a game with time-homogeneous rules played on the translation-invariant real line—how could any finite-time expenditure be justified (in furtherance of claiming an ultimate finite reward) when the future from $(t,X(t))$ is indistinguishable from the time-zero prospect?—is thus resolved: the space of equilibria is invariant under $\mathbb{R}$-shift, but symmetry breaks for the elements, with each distinguishing a zone in the real line where the true battle will take place.
Safe to say, we need only study the solution quadruple for a single value of $x$, with the most convenient choice being $x=1$. So the next result captures an important aspect of all solutions’ behaviour.
Proposition 1.5.
For $\rho\in(0,\infty)$, the functions $f_{\rho}(1,u)$, $g_{\rho}(1,u)$, $a_{\rho}(1,u)$ and $b_{\rho}(1,u)$ take the form $u^{\zeta}\,e^{-2u}\,\Theta(1)$ as $u\to\infty$. The exponent $\zeta$ is determined by the function via
$\zeta_{f}=\frac{1+\rho}{2\rho^{2}}\,\,\,,\,\,\,\zeta_{g}=\frac{1-\rho}{2\rho^{2}}\,\,\,,\,\,\,\zeta_{a}=\frac{1+\rho}{2\rho^{2}}-1\,\,\,\,\textrm{and}\,\,\,\,\zeta_{b}=\frac{1-\rho}{2\rho^{2}}-1\,.$
As $u\to$-$\infty$, the functions’ form is $|u|^{\zeta}e^{-2|u|}\Theta(1)$ after interchanges of $\zeta_{f}$ and $\zeta_{g}$ and of $\zeta_{a}$ and $\zeta_{b}$.
Figure $1.$ The curves $a=a_{1}(1,\cdot)$ and $f=f_{1}(1,\cdot)$, as specified by Definitions 1.2 and 1.4; $b$ and $n$ are given by reflecting in the vertical axis. Maxine’s stake profile $a$ takes maximum value $0.57$ at $x=0.25$ to two decimal places.
When $x=1$, a bounded neighbourhood of the origin may be viewed as the site of a battlefield, where players stake at unit order. (As Figure 1.1 depicts in the case $\rho=1$, a player spends most as she begins to lead.) All four functions, including the stake profiles $a_{\rho}(1,\cdot)$ and $b_{\rho}(1,\cdot)$, decay rapidly, as $e^{-2|u|}$, at high distances from the origin. There is a more modest but clear asymmetry in the rate of decay, manifest in the values of the power-law exponents. When $u\gg 0$, the presumptively leading player, Maxine, is staking at normalized rate $u^{\zeta_{a}}$, above the analogous level of $u^{\zeta_{b}}$ for Mina; and this circumstance is swapped in the opposite regime $u\ll 0$. As $\zeta_{a}>\zeta_{b}$, so $\zeta_{f}>\zeta_{g}$: $1\gg m^{\prime}\gg-n^{\prime}>0$ when $u\gg 0$; integrating on $u,\infty)$, the leading player’s shortfall in expectation relative to her winning terminal receipt is seen to exceed the opponent’s excess over her losing terminal receipt. This imbalance reflects the effort of expenditure that the leading player must exert—small in an absolute sense, but large relative to the opponent’s—in order to convert a likely victory.
The rich yet explicit structure of solutions in Theorem 1.3 appears to mark $\rho$-Brownian Boost as an outlier among analysed non-zero-sum stochastic differential games. Zero-sum examples with explicit solutions include the stochastic linear-quadratic regulator problem [21, Example $III.8.1$], and fair-coin tug-of-war [41], whose infinity-harmonic value functions [3] take explicit forms in certain cases, particularly in two dimensions.
The stake pairs in Definition 1.4 arising from the ODE solutions in Theorem 1.3 offer a classification of all time-homogeneous Markov perfect Nash equilibria in $\mathrm{BB}(\rho)$. We will utilize discrete counterparts of Brownian Boost, in a fine-mesh high-noise limit, to substantiate this assertion rigorously, in Theorem 1.18. The discrete games are akin to random-turn tug-of-war games that have been considered since the 1980s in the economics literature of dynamic contests. At the same time, the analytic solutions for Brownian Boost form a point of contact with a parallel but oddly disjoint tug-of-war research vein in probability and PDE that dates from the 2000s. In this way, we hope that Brownian Boost may offer a bridge between the two tug-of-war literatures.
An overview of the two literatures will provide context for discussing the prelimiting discrete games. In the next subsection, we offer one, indicating at the end the structure of the rest of the introduction.
@@ 1.2. Tug of war, in economics and mathematics
In 2009, Peres, Schramm, Sheffield and Wilson [41] studied a class of random-turn games, which they also named tug-of-war. Played on a discrete graph $G=(V,E)$ with boundary $B$, or in a domain $D$ in Euclidean space, the game begins with a counter located at a vertex in $V$ or at an interior point of $D$. At each turn, a fair coin is flipped and the turn victor moves the counter to a location of his choosing: an adjacent vertex in the discrete setting; and, in the continuous one, a point in $D$ at distance at most $\epsilon$ away, where $\epsilon>0$ is a parameter fixed for the game. On the boundary $B$ or $\partial D$ is specified a real-valued payment function $f$. The game ends when the counter arrives in the boundary with a payment from one player to the other given by the evaluation of $f$ at the terminal counter location. In the discrete setting, the game value $h(v)$ expressed as a function of starting location $v$ is the extension of $f$ that satisfies $h(v)=\big(\max_{u\sim v}h(u)+\min_{u\sim v}h(u)\big)/2$, the minimum and maximum over neighbours reflecting the choices made when playing from $v$. The equation is an $\infty$-version of the mean value property in which only the two extremes contribute to the average. In the Euclidean setting, the infinity-harmonic extension of $f$ to $D$ is the viscosity solution $h:D\to\mathbb{R}$ of the infinity Laplace equation $\sum_{i,j}\partial_{x_{i}}h\,\partial_{x_{i}x_{j}}h\,\partial_{x_{j}}h=0$ subject to $h\big|_{B}=f$, whose second derivative in the gradient direction vanishes. In [41], it is proved the value of tug-of-war (as these authors named the game) played on $D$ converges in the low-$\epsilon$ limit to this extension.
@@ 1.2.1. The economics vein
One rule to model a single round in player-funded tug-of-war is a Tullock contest [44]. This is a single-stage game in which player $A$ stakes $x\in[0,\infty)$ and player $B$, $y\in[0,\infty)$, the contest won by $A$ with probability $\tfrac{x^{\rho}}{x^{\rho}+y^{\rho}}$, where $\rho\in(0,\infty)$ is now called the Tullock exponent. When $\rho\to\infty$, all-pay auctions are obtained, in which the higher staking player wins. Player-funded tug-of-war, including the role of battlefields and discouragement, has been studied [32, 1, 33] on finite integer intervals with the all-pay auction rule used to decide turn victor and in variants [25, 24] where a firm is composed of several individuals who are responsible for different payments. The player-funded game has been studied with the majoritarian objective in which the patent is awarded to the firm who first achieves a certain number of turn victories, as a model of the premise that early expenditure is decisive, in [31]; with intermediate prizes [34]; and with discounting [23] viewed as a dissipator of momentum for the leading player. Two phases of play—no site revisits, then tug of war—occur in a more general graphical framework studied in [18].
A separate thread in the economics literature concerns stake-governed tug-of-war where, rather than pay from their own savings, players finance their stakes from a budget allocated to them as part of the game design. See [30] for an analysis with the majoritarian objective, and [29] for finite integer intervals. In [22], a leisurely or lazy version of the game is studied on a class of trees, with connections drawn to constant-bias tug-of-war.

### v4

@@ 1.1.2. Some signposts for Brownian Boost results
In the preface, we stated two premises concerning Brownian Boost: that it is unplayable because of the translation invariance of its real-line gameboard; and that any incentive asymmetry disrupts non-trivial solution structure. Considering this pair of notions may offer a useful route towards understanding Brownian Boost and its mathematical structure. Later in the introduction, we will introduce the discrete Trail of Lost Pennies games and explicitly classify their TINE; and then scale them to find a counterpart classification for Brownian Boost. Doing so will naturally involve some elaboration in notation and statement. Here, we offer some signposts to the form of the Brownian Boost equilibria we will exhibit. While what we indicate is merely a fragment of the solution structure, this piece of the picture is enough to permit us to evaluate the validity of the two premises.
Our analytic presentation of these equilibria is governed by two coupled ODE, the $\mathrm{BB}(\rho)$-ODE pair (1.13), whose solutions describe (up to sign) the spatial derivatives of the players’ value functions; the corresponding equilibrium stake profiles (or TINE) and the induced counter dynamics are then expressed as simple functionals of this solution pair. Here we describe precisely only one aspect, the dynamics, which admits a simple description.
The set of TINE in $\mathrm{BB}(\rho)$ will be classified when $\rho\in(0,1]$. For each such $\rho$, this collection is a one-dimensional space, invariant under real shifts, and indexed by a ‘battlefield’ value in $\mathbb{R}$. From the equilibrium $(a,b):\mathbb{R}\to(0,\infty)$ with battlefield value zero, all other TINE are obtained in the form $\big(a(\bullet-v),b(\bullet-v)\big)$ as the battlefield value $v$ varies over $\mathbb{R}$. We defer recording the form of $(a,b)$, and instead indicate the gameplay that results when the players follow this zero-indexed TINE. This gameplay $X:[0,\infty)\to\mathbb{R}$ solves
${\rm d}X_{t}=R_{\rho}(X_{t})\,{\rm d}t+{\rm d}W_{t}\,,\,\,\,\textrm{with $W_{t}$ standard Brownian motion}\,$, (1.1)
an SDE in whose drift term $R_{\rho}:\mathbb{R}\to(-1,1)$ is given by $R_{\rho}(u)=\frac{1-J(u)}{1+J(u)}$ where $J:\mathbb{R}\to(0,\infty)$ is a decreasing bijection that is the unique solution of the ODE
$\frac{{\rm d}J(u)}{{\rm d}u}\,=\,-8\rho^{2}\frac{J(u)^{2}}{(1+J(u))^{2}}$
with $J(0)=1$. The function $R_{\rho}$ has qualitative similarities to the $(2/\pi)$-multiple of the arctangent: it odd, smooth, and satisfies $R_{\rho}(u)\to\pm 1$ as $u\to\pm\infty$.
Brownian Boost is a playable game when $\rho\in(0,1]$, in the sense that solutions exist for a standard notion of equilibrium. The first premise advanced in the preface is false because the translation-invariance of the real line is manifest in the space of TINE; this symmetry breaks for any given element of this set. Indeed, the deferred explicit form of the zero-indexed equilibrium $(a,b):\mathbb{R}\to(0,\infty)^{2}$ comprises two functions of unit order in a compact neighbourhood of the origin, with exponential decay at higher distances. When this TINE is played, both players vie for control of the counter near the origin, rapidly curtailing stakes at further distances. The origin (and its locale) is a battlefield where the outcome is typically decided.
What of the second premise, that no asymmetry in incentive is tolerated in Brownian Boost? The solution $J$ of the above initial value problem is straightforwardly seen to satisfy J(u)=1/J(u)J(-u)=1/J(u); we will obtain this equality in the proof of Proposition 5.4(1,2). Consequently, the zero-indexed TINE gameplay SDE has a drift term that is an odd function of spatial location (as we stated above). This gameplay symmetry is in fact incompatible with any asymmetry in terminal receipt. Which is to say: the value $\lambda$ of Mina’s terminal receipt in the event of her victory is equal to one. This conclusion holds for the zero-battlefield TINE for $\rho\in(0,1]$; and it holds for any other TINE by translation invariance of the TINE set. So the second premise is validated for Brownian Boost: no asymmetry in incentive is permissible.
With these signposts indicated, we are ready to turn to specify the form and solution of the Trail of Lost Pennies games that underpin our Brownian Boost analysis. The game studied in [24] is the special case $\mathrm{TLP}(1,1)$ of a two-parameter family $\mathrm{TLP}(\kappa,\rho)$ that we consider. Our study of these games was prompted by a referee of [24] who asked whether mixing the game in question with random-walk noise changes the dynamics. There are two independent directions in the broadening we consider: a noise parameter $\kappa\in(0,1)$ that mixes stake-governed motion with fair coin flips, per the referee’s suggestion; and an exponent $\rho\in(0,\infty)$ that specifies the Tullock contest on which each turn is modelled. In this way, the one-dimensional collection of Brownian Boost games, parameterised by $\rho$, is rigorously analysed as a regularization of the discrete games. This is achieved by taking a fine-mesh, high-noise limit $\kappa\searrow 0$.
These discrete games are akin to random-turn tug-of-war games that have been considered since the 1980s in the economics literature of dynamic contests; indeed, beyond their role in interpreting Brownian Boost, the Trail of Lost Pennies games form a rich source for examining prevalent themes in that literature. It may thus be instructive to first review this vein of research; and any such review for mathematicians naturally raises the prospect of comparison with a parallel but oddly disjoint tug-of-war research vein in probability and PDE that dates from the 2000s.
In the next section, we indeed offer a review of these two literatures. Three further sections then complete the introduction. The first addresses the setup and solution of the discrete games; the second, how they scale to Brownian Boost, and the resulting classification of equilibria for the continuum game.
While the discouragement effect may be said to be absolute in Brownian Boost—all TINE have $\lambda=1$—the counterpart question for the discrete games has a nuanced answer which develops that found for $\mathrm{TLP}(1,1)$ in [24]: asymmetric perturbation is highly fragile but not infinitely so, with the degree of this fragility depending on the parameter pair $(\kappa,\rho)$ in non-trivial and in some cases very surprising ways. In the final section of the introduction, we examine how the explicit forms for equilibria in the discrete and continuous games cast light on themes in the economics literature such as discouragement, explaining how numerical evidence supports such conclusions about fragility under incentive asymmetry.
@@ 1.2. Tug of war, in economics and mathematics
In 2009, Peres, Schramm, Sheffield and Wilson [43] studied a class of random-turn games, which they also named tug-of-war. Played on a discrete graph $G=(V,E)$ with boundary $B$, or in a domain $D$ in Euclidean space, the game begins with a counter located at a vertex in $V$ or at an interior point of $D$. At each turn, a fair coin is flipped and the turn victor moves the counter to a location of his choosing: an adjacent vertex in the discrete setting; and, in the continuous one, a point in $D$ at distance at most $\epsilon$ away, where $\epsilon>0$ is a parameter fixed for the game. On the boundary $B$ or $\partial D$ is specified a real-valued payment function $f$. The game ends when the counter arrives in the boundary with a payment from one player to the other given by the evaluation of $f$ at the terminal counter location. In the discrete setting, the game value $h(v)$ expressed as a function of starting location $v$ is the extension of $f$ that satisfies $h(v)=\big(\max_{u\sim v}h(u)+\min_{u\sim v}h(u)\big)/2$, the minimum and maximum over neighbours reflecting the choices made when playing from $v$. The equation is an $\infty$-version of the mean value property in which only the two extremes contribute to the average. In the Euclidean setting, the infinity-harmonic extension of $f$ to $D$ is the viscosity solution $h:D\to\mathbb{R}$ of the infinity Laplace equation $\sum_{i,j}\partial_{x_{i}}h\,\partial_{x_{i}x_{j}}h\,\partial_{x_{j}}h=0$ subject to $h\big|_{B}=f$, whose second derivative in the gradient direction vanishes. In [43], it is proved the value of tug-of-war played on $D$ converges in the low-$\epsilon$ limit to this extension.
@@ 1.2.1. The economics vein
One rule to model a single round in player-funded tug-of-war is a Tullock contest [47]. As we discussed in the preface, this is a single-stage game in which player $A$ stakes $x\in[0,\infty)$ and player $B$, $y\in[0,\infty)$, the contest won by $A$ with probability $\tfrac{x^{\rho}}{x^{\rho}+y^{\rho}}$, where $\rho\in(0,\infty)$ is the Tullock exponent. This family of win probabilities (or contest success functions) is a natural class of expressions that are homogeneous of degree zero. Jia [27] has shown that the contest success function must in fact take the Tullock form when finitely many players compete in a contest in which the victor is the player with the highest realized performance, defined as the product of her stake and an independent random factor with a law shared across players, provided that the vector of win probabilities conditional on the stake vector satisfies Luce’s choice axiom, which encodes the independence of irrelevant alternatives. From an axiomatic standpoint, Skaperdas [45] shows that basic structural requirements—scale invariance, monotonicity, and a contest-level independence-of-irrelevant-alternatives principle—again single out the Tullock form.
When $\rho\to\infty$ in the Tullock contest, all-pay auctions are obtained, in which the higher staking player wins. Player-funded tug-of-war, including the role of battlefields and discouragement, has been studied [32, 1, 33] on finite integer intervals with the all-pay auction rule used to decide turn victor and in variants [23, 22] where a firm is composed of several individuals who are responsible for different payments. The player-funded game has been studied with the majoritarian objective in which the patent is awarded to the firm who first achieves a certain number of turn victories, as a model of the premise that early expenditure is decisive, in [30]; with intermediate prizes [34]; and with discounting [21] viewed as a dissipator of momentum for the leading player. Two phases of play—no site revisits, then tug of war—occur in a more general graphical framework studied in [17].
A separate thread in the economics literature concerns a form of stake-governed tug-of-war where, rather than pay from their own savings, players finance their stakes from a budget allocated to them as part of the game design. See [29] for an analysis with the majoritarian objective, and [28] for finite integer intervals. In [20], a leisurely or lazy version of the game is studied on a class of trees, with connections drawn to constant-bias tug-of-war.